OpenStudy (anonymous):

Solve using Linear Combination 2e-3f+-9 e+3f=18 Will medal/fan for best correct answer

rebeccaxhawaii (rebeccaxhawaii):

both -3f and +3f = 0

OpenStudy (ask_if_im_a_ghost):

Because in both equations the 3f's have opposite signs, you can add the two equations, giving you\[2e+e = 18-9\]Then you solve for e. After obtaining e, you plug it into any of the two equations then solve for f.

satellite73 (satellite73):

\[ 2e-3f=-9\\ e+3f=18\]

rebeccaxhawaii (rebeccaxhawaii):

the f's cancel out. letting you solve for e.

rebeccaxhawaii (rebeccaxhawaii):

 therefore 2e+e=-9+18